OpenStudy (aaronandyson):

The sum S of n successive odd number starting from 3 is given by the relation. S = n(n+2).Determine n,If the sum is 168

OpenStudy (aaronandyson):

@Sepeario

OpenStudy (sepeario):

Ok so substitute values for S first.

OpenStudy (aaronandyson):

S = n(n+2) 168 = n(n+2) 168 = n^2 + 2n n^2 + 2n - 168 = 0

OpenStudy (sepeario):

Okay and can you factorise that now?

OpenStudy (aaronandyson):

No :(

OpenStudy (sepeario):

give me two seconds plz.

OpenStudy (sepeario):

The factorised form for this quadratic is (n-12)(n+14)

OpenStudy (sepeario):

Hence the solutions: -12 and 14.

OpenStudy (aaronandyson):

Thanks Can u help me more?

OpenStudy (sepeario):

idk how much more?